OpenStudy (anonymous):

If the first step in the solution of the equation -9 + x = 5x - 7 is "subtract x," then what should the next step be?

OpenStudy (anonymous):

add 7 makes more sense to try to solve it than to explain the steps

OpenStudy (anonymous):

\[-9 + x = 5x - 7 \] subtract \(x\)\[-9=4x-7\] then you have \(x\) on the left, you want the number on the right, so add 7 get \[-2=4x\]

OpenStudy (anonymous):

finish by dividing by 4 to get \[-\frac{1}{2}=x\]